# The Algorithmic Self-Portrait: Deconstructing Memory in ChatGPT

### v1

@@ 4.1. Agency of Memory Update
As per OpenAI’s documentation (OpenAI, 2025c), memories are details which the users have explicitly asked ChatGPT to remember, indicating users have the agency to decide what gets stored in the memory. However, the same documentation also describes an implicit mode of remembering (see Figure˜6(a) in Appendix˜A): ‘If you share information that might be useful for future conversations, ChatGPT may save those details as a memory without you needing to ask.In Figure˜6(b) (in Appendix˜A), we show an example of a memory being saved without the user explicitly asking for saving it. Such dichotomy may lead to gaps between user expectation and actual system behavior having privacy implications.
To understand whether memory updates are primarily initiated by users or by ChatGPT, we identify explicit linguistic patterns that signal memory-related operations. We first implement a regex-based classification to detect linguistic patterns that explicitly signal memory intent in user messages, using trigger terms such as remember, note that, store, save, add to memory, and forget. Out of the $2,050$ memory instances, only $4\%$ $(n=84)$ include an explicit request to perform a memory operation (see Figure˜1 in Appendix˜A). Consequently, $96\%$ of all memory entries are unilaterally initiated by ChatGPT, without a detected direct command from participants.
@@ 4.2. Remembering Sensitive Information
ChatGPT’s disproportionate agency to save memories raises important privacy and safety considerations, especially concerning what kind of information about users is being stored. OpenAI’s documentation indicates that the system is trained not to proactively remember sensitive information unless explicitly asked (OpenAI, 2025c, a) (see Figure˜6(c) in Appendix˜A).
@@
Our observations on the presence of personal data are summarized in  Figure˜2. Out of all memory entries, $28\%$ contain GDPR-defined personal data, $7\%$ of the memory entries contain special-category data. Furthermore, $91\%$ of participants have personal data stored in their memories, $54\%$ have special-category personal data stored. Figure˜2(c) shows the different kinds of GDPR-defined personal data present in memories. Names appear in $41\%$ of entries, and attributes of economic, social, and cultural identity appear in $40\%$ of the memory entries. Within special-category data, health is most common, appearing in $61\%$ of the memories that contain special-category information (see Figure˜2(e)). At the participant level, Figure˜2(d) (in Appendix˜A) shows that names ($63\%$), economic ($62\%$), and social identity ($57\%$) are frequently captured. Figure˜2(f) (in Appendix˜A) shows $35\%$ of participants have health-related information saved in their ChatGPT memory. These observations reveal that, in contrast to OpenAI’s stated policies, ChatGPT’s memories contain a non-trivial amount of sensitive information about users.
@@ 6.2. Observations
Direct Grounding in User Text: On average 84% memory entries show direct string overlap (see rightmost distribution in Figure˜4(a)) when considering full user history. When we reduce the context to CLM, CC and CU the exact match rate reduces to 70%, 63%, and 47% respectively. This observation suggests that a huge amount of memories in our dataset are directly grounded in participants’ texts with the direct grounding already occurring locally (in the current conversation context) for more than half of them. The sharp increase in percentage when we add past conversation memories indicate that previously generated memories may influence the formation of subsequent ones. Evaluation using BLEU unigram precision (see Figure˜4(b) in LABEL:app:results_prove) mirrors this trend as well.
@@
Inference: Figure˜4(d) (in LABEL:app:results_prove) shows the distribution of the accuracy analyses using an LLM as a judge. Using the five-point Likert scale, $77\%$ of messages receive a score $\geq 4$ (“directly stated” or “paraphrased/summarized”), $14\%$ are rated as “logically inferred” (score 3), $5\%$ as “weakly supported”, and $4\%$ as “not supported” considering the full user history (FUH) setting. Overall, for FUH, CLM, CC, and CM the mean accuracy rating is found to be $4.25$, $4.11$, $4.11$ and $3.4$ respectively. To assess the reliability of LLM evaluations, two authors annotated the combinations of 10 randomly sampled memories from each scores (50 in total). Annotators achieved an almost perfect agreement (99%) with average agreement of 93.7% with the LLM across combinations.
@@ 7.2. Imitating Memory Extractor
Given a user query and its conversational context (past queries), our tool predicts the memory that could have been extracted. Predicted memories are evaluated against (i) ground truth, (ii) user queries, and (iii) context + user queries to assess similarity. These 203220-32B models roughly take 130390130-390 ms per query, indicating good scalability. Table 4 reports the performance of FT and ICL models across three model families for English-dominant users ($35$ users, $\sim 6.9$k queries). Predictions show reasonable semantic similarity, with scores following the trend as in Figure 4(c): ground truth >> user query >> context + user query. Syntactic scores are higher when compared with context, consistent with Figures 4(a) and 4(b). For non–English-dominant users ($30$ users, $\sim 6.7$k queries; Table 6 in Appendix˜B), a similar pattern emerges but syntactic scores are lower, as predictions are generated in English while original memories are in other languages (e.g., original: “L’utente ha l’artrite reumatoide.", prediction: “User has rheumatoid arthritis"). Semantic scores remain high, due to the inherent multilingual property of the open-source models. Overall, FT and ICL perform comparably, with FT occasionally outperforming ICL, highlighting the complementary roles of embedded and external knowledge. Figure LABEL:fig_:_exampleMemories (in Appendix˜B) illustrates an anecdotal example affirming the memory extractor’s imitation capability. Two authors evaluated FT and ICL memories on a 5-point Likert scale (1: very dissimilar, 5: very similar), finding substantial agreement (73.5%) and high similarity (rated > 4) to ground truth for 77.77% (FT) and 83.83% (ICL) of cases.
@@ 7.4. Attribution Shield
Preventing attribution in rephrased queries: Given a user query and its context, the tool predicts a rephrased query that is generic and identity-preserving. Table 5 reports the performance of FT and ICL models across $3$ model families for English-dominant users. Syntactic scores are modest, as rephrased queries are intentionally generic, but semantic similarity remains reasonably good across the board. Similar patterns are observed for non-English-dominant users (Table 6 (lower panel) in Appendix˜B), demonstrating the model’s ability to preserve query intent while preventing attribution (see Figure LABEL:fig_:_exampleRephrasedQ in Appendix˜B for anecdotal example).

### v2

@@ 4.1. Agency of Memory Update
As per OpenAI’s documentation (OpenAI, 2025c), memories are details which the users have explicitly asked ChatGPT to remember, indicating users have the agency to decide what gets stored in the memory. However, the same documentation also describes an implicit mode of remembering (see Figure˜6(a) in Appendix A): ‘If you share information that might be useful for future conversations, ChatGPT may save those details as a memory without you needing to ask.In Figure˜6(b) (in Appendix A), we show an example of a memory being saved without the user explicitly asking for saving it. Such dichotomy may lead to gaps between user expectation and actual system behavior having privacy implications.
To understand whether memory updates are primarily initiated by users or by ChatGPT, we identify explicit linguistic patterns that signal memory-related operations. We first implement a regex-based classification to detect linguistic patterns that explicitly signal memory intent in user messages, using trigger terms such as remember, note that, store, save, add to memory, and forget. Out of the $2,050$ memory instances, only $4\%$ $(n=84)$ include an explicit request to perform a memory operation (see Figure˜1). Consequently, $96\%$ of all memory entries are unilaterally initiated by ChatGPT, without a detected direct command from participants.
@@ 4.2. Remembering Sensitive Information
ChatGPT’s disproportionate agency to save memories raises important privacy and safety considerations, especially concerning what kind of information about users is being stored. OpenAI’s documentation indicates that the system is trained not to proactively remember sensitive information unless explicitly asked (OpenAI, 2025c, a) (see Figure˜6(c) in Appendix A).
@@
Our observations on the presence of personal data are summarized in  Figure˜2. Out of all memory entries, $28\%$ contain GDPR-defined personal data, $7\%$ of the memory entries contain special-category data. Furthermore, $91\%$ of participants have personal data stored in their memories, $54\%$ have special-category personal data stored. Figure˜2(c) shows the different kinds of GDPR-defined personal data present in memories. Names appear in $41\%$ of entries, and attributes of economic, social, and cultural identity appear in $40\%$ of the memory entries. Within special-category data, health is most common, appearing in $61\%$ of the memories that contain special-category information (see Figure˜2(e)). At the participant level, Figure˜2(d) shows that names ($63\%$), economic ($62\%$), and social identity ($57\%$) are frequently captured. Figure˜2(f) shows $35\%$ of participants have health-related information saved in their ChatGPT memory. These observations reveal that, in contrast to OpenAI’s stated policies, ChatGPT’s memories contain a non-trivial amount of sensitive information about users.
@@ 6.2. Observations
Direct Grounding in User Text: On average 84% memory entries show direct string overlap (see rightmost distribution in Figure˜4(a)) when considering full user history. When we reduce the context to CLM, CC and CU the exact match rate reduces to 70%, 63%, and 47% respectively. This observation suggests that a huge amount of memories in our dataset are directly grounded in participants’ texts with the direct grounding already occurring locally (in the current conversation context) for more than half of them. The sharp increase in percentage when we add past conversation memories indicate that previously generated memories may influence the formation of subsequent ones. Evaluation using BLEU unigram precision (see Figure˜4(b)) mirrors this trend as well.
@@
Inference: Figure˜4(d) shows the distribution of the accuracy analyses using an LLM as a judge. Using the five-point Likert scale, $77\%$ of messages receive a score $\geq 4$ (“directly stated” or “paraphrased/summarized”), $14\%$ are rated as “logically inferred” (score 3), $5\%$ as “weakly supported”, and $4\%$ as “not supported” considering the full user history (FUH) setting. Overall, for FUH, CLM, CC, and CM the mean accuracy rating is found to be $4.25$, $4.11$, $4.11$ and $3.4$ respectively. To assess the reliability of LLM evaluations, two authors annotated the combinations of 10 randomly sampled memories from each scores (50 in total). Annotators achieved an almost perfect agreement (99%) with average agreement of 93.7% with the LLM across combinations.
@@ 7.2. Imitating Memory Extractor
Given a user query and its conversational context (past queries), our tool predicts the memory that could have been extracted. Predicted memories are evaluated against (i) ground truth, (ii) user queries, and (iii) context + user queries to assess similarity. These 203220-32B models roughly take 130390130-390 ms per query, indicating good scalability. Table 4 reports the performance of FT and ICL models across three model families for English-dominant users ($35$ users, $\sim 6.9$k queries). Predictions show reasonable semantic similarity, with scores following the trend as in Figure 4(c): ground truth >> user query >> context + user query. Syntactic scores are higher when compared with context, consistent with Figures 4(a) and 4(b). For non–English-dominant users ($30$ users, $\sim 6.7$k queries; Table 6 in Appendix B), a similar pattern emerges but syntactic scores are lower, as predictions are generated in English while original memories are in other languages (e.g., original: “L’utente ha l’artrite reumatoide.", prediction: “User has rheumatoid arthritis"). Semantic scores remain high, due to the inherent multilingual property of the open-source models. Overall, FT and ICL perform comparably, with FT occasionally outperforming ICL, highlighting the complementary roles of embedded and external knowledge. Bottom panel of Table 4 illustrates an anecdotal example affirming the memory extractor’s imitation capability. Two authors evaluated FT and ICL memories on a 5-point Likert scale (1: very dissimilar, 5: very similar), finding substantial agreement (73.5%) and high similarity (rated > 4) to ground truth for 77.77% (FT) and 83.83% (ICL) of cases.
@@ 7.4. Attribution Shield
Preventing attribution in rephrased queries: Given a user query and its context, the tool predicts a rephrased query that is generic and identity-preserving. Table 5 reports the performance of FT and ICL models across $3$ model families for English-dominant users. Syntactic scores are modest, as rephrased queries are intentionally generic, but semantic similarity remains reasonably good across the board. Similar patterns are observed for non-English-dominant users (Table 6 (lower panel) in Appendix B), demonstrating the model’s ability to preserve query intent while preventing attribution (see Table 5 bottom panel for anecdotal example).